OpenStudy (anonymous):

Snookers Lumber can convert logs into either lumber or plywood. In a given day, the mill turns out 3 times as many units of plywood as lumber. it makes a profit of $30 on a unit of lumber and $50 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of $18360?

OpenStudy (anonymous):

The first equation: Profit is the number of units times the money made per unit. This is calculated for both Lumber and Plywood. Units_of_Lumber($30) + Units_of_Plywood($50) = $18,360 30L + 50P = 18360 The second equation: The problem states that there is three times the amount of plywood produce than lumber. Units_of_Plywood = 3 * Units_of_Lumber P = 3L Substitute 3L for P. 30L + 50(3L) = 18360 Solve. 30L + 150L = 18360 180L = 18360 L = 102 Units of Lumber P = 3L = 306 Units of Plywood